Quadratic equation
4. The product of two consecutive numbers is 3306. What are the numbers?
Thanks!!!

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

The product of two consecutive numbers is 3306. What are the numbers?
:
Two consecutive numbers, x, (x+1)
:
The product:
x * (x+1) = 3306
:
x^2 + x = 3306
:
A quadratic equation
x^2 + x - 3306 = 0
:
Find the square root of 3306 = 57.497.. we know the x value will be close to that
(x + 58)(x - 57) = 0
Two solutions
x = 57, 58 are the two numbers
and
x = -58, -57 are also be a solution

Answer by Alan3354(69443)

A quicker method:
sqrt(3306) = ~ 57.49
The integers are 57 and 58, and -57 and -58